Session: May 1999 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Sat May 29 19:02:57 METDST 1999) Success

Torun (K. Borkowski at Sun May 30 15:47:47 METDST 1999) Success
The schedule did not allow enough time for the telescope pass to a new source in a few cases.

Onsala (A. Polatidis at Sun May 30 17:07:23 METDST 1999) Minor failures
No problems with the observations itself. The problem with the recording quality persists, and it is worse in a double speed experiment such as this. We made an attempt to recontour the heads after the end of this experiment and we have slightly improved the quality.

Westerbork (Tony Foley at Mon May 31 15:50:29 METDST 1999) Minor failures
Used 12 telescopes in tied-array until end, but one of them (RTA) got stuck from 14:05-15:00 UT. VC06 would not respond (not in use anyway). Parity error checks using thin tape procedures gave silly answers, but playback at 135ips after the experiment looked acceptable .

Noto (C.Trigilio at Mon May 31 16:52:19 METDST 1999) Minor failures
First two scans probably affacted by a formatter error of -2 sec. Formatter reset during 3 scan (lost). No problem later on.

Effelsberg (A. Kraus at Tue Jun 1 09:29:57 METDST 1999) Minor failures
Few Scans have been lost due to insufficient slewing time.

Jodrell (Alastair Gunn at Tue Jun 1 15:55:39 METDST 1999) Success
No known problems.

Cambridge (Alastair Gunn at Tue Jun 1 15:56:08 METDST 1999) Minor failures
First two scans have bad data due to wrong LO configuration. Tape failed to start for first scan anyway. No other problems.

Calibration (Jean\-Francois Desmurs at Fri Jun 18 19:32:53 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may99 directory under the name: es023b.antab. Read the comments in the file to get further information.